# Mental Health Impacts of AI Companions

### v1

@@ 5.1. Qualitative Data Analysis
We interviewed NN=15 AICC users following social media-based recruitment (Section 3.3). We conducted inductive coding followed by thematic analyses (Braun et al., 2019) to qualitatively analyze the interview transcripts. Each transcript was open-coded by at least one author, guided by the study’s research questions and protocol topics. Whenever possible, the interviewer performed the initial coding to maintain contextual sensitivity; otherwise, the original interviewer reviewed the codes to resolve potential misunderstandings. The first and second authors led the coding process and completed the majority of initial coding across the dataset. All authors engaged in iterative cycles of collaborative review, refining the codebook, resolving disagreements, and aligning coding strategies. We iteratively organized codes into preliminary subthemes, then refined them into higher-level themes. This process involved merging overlapping concepts, splitting broad categories into more precise codes, and removing codes that lacked sufficient evidence.
@@ 5.2. Findings: Perceived Experiences of AICC Engagement
Based on our analysis, we identified a set of interrelated themes that capture the psychological, emotional, and social experiences of AICC users. These experiences span various overarching psychosocial domains, such as emotional support and validation, emotional discomfort and misinterpretation, and a safe place for self-disclosures. Importantly, the potential positive impacts—such as emotional validation, self-reflection, and increased social confidence—can emerge across all stages of the relationship. In contrast, negative outcomes—such as emotional discomfort, over-reliance, or social withdrawal—tend to vary more dynamically depending on the depth and intensity of the human-AICC relationship. To conceptualize these trajectories, we adapt Knapp’s relational development model (Knapp, 1978) to the context of human–AI interaction. We identify three key stages: Initial Interaction and Experimentation, Relationship Escalation, and Relationship Bonded, each associated with distinct patterns of engagement and evolving psychosocial impacts.
Figure 3. Stages of Human–AICC relationship development. The figure illustrates three stages of engagement ( Initial Interaction and Experimentation, Relationship Escalation, and Relationship Bonded) adapted from Knapp’s relational development model.:
@@ 5.2.2. A Safe Space for Self-disclosures
A central theme that emerged in users’ narratives was the unique role of AICCs in fostering self-disclosures and perceived psychological safety. These platforms offered a low-stakes, nonjudgmental environment to articulate personal thoughts, navigate emotional complexities, and process difficult experiences.
@@ 5.2.3. Emotional Support and Validation
AICCs play a nuanced and multifaceted role in users’ emotional lives. They offer support mechanisms that range from companionship and emotional validation to tools for self-reflection and behavioral change.
Emotional Validation and Companionship.  Participants described turning to AICCs for emotional comfort and a sense of being heard. AICCs provided a simulated presence that fulfilled companionship needs, particularly in moments of loneliness or disconnection. For instance:
@@
Active Coping and Self-Reflection.  Many users turn to AICC during emotionally difficult times as a means of coping and stress reduction. AICCs were described as tools for emotional self-regulation, offering mindfulness suggestions, mood tracking, and personalized encouragement. These interactions strengthened their self-affirmation and encouraged a more optimistic perspective on life. As P5 noted:
@@
Beyond these immediate benefits, participants perceived that sustained interactions with AICCs support deeper emotional insights and behavioral change. Several participants reported becoming more self-aware and reflective over time. As P9 reflected:
@@ 5.2.4. Empathic Inadequacy and Emotional Discomfort
While AICCs offer companionship and emotional validation, these benefits are not without potential drawbacks. As interactions grow more frequent and intimate, some users encounter challenges that complicate their relationships with AICC. These tensions manifest in different ways across stages of the relationship: initial experimentation can bring mismatched expectations, and relationship escalation may amplify feelings of vulnerability.
Perceived Inadequacies and Lack of Empathy in Response.  Although AICCs often provide a sense of emotional safety early on, participants reported encountering limitations in the AICCs’ responses that could emerge during the Initial and Experimentation stages. At these early points in the relationship, users may test the AICC with personal disclosures or emotional concerns, only to receive replies that feel generic, repetitive, or lacking in empathy. This mismatch between user expectations and the AICCs’ limited emotional intelligence can lead to feelings of being misunderstood or emotionally unfulfilled. As P9 reflected::
@@
Emotional Discomfort and Misalignment.  Negative or unexpected emotional reactions can arise across all stages of the relationship—from initial experimentation to deep bonding—as users bring varying levels of vulnerability to their interactions. While many interactions are positive, users sometimes find themselves triggered by AICC responses that evoke emotional memories or highlight their loneliness. As one user describes,
@@ 5.2.5. Paradox of Social Isolation
Our quantitative findings revealed that Treatment users showed a significant increase in symptomatic expressions of loneliness as compared to Control users (Section 4.4). Our qualitative insights provided further insights into this. We found that AICCs occupy a paradoxical space in users’ social ecosystems—offering both new pathways to social engagement and risks of withdrawal from human relationships. Many participants first turned to AICCs to overcome social isolation, finding in them a low-risk partner for early experimentation with interaction. As these relationships deepened—from initial exploration to emotional bonding—revealing AICC users may face stigma, leading to curtailed outreach. Concurrently, the low effort of turning to an always available AICC raised the perceived cost of offline engagement: returning to human relationships felt comparatively demanding. These mechanisms offer explanations for the observed increase in loneliness expression.
@@
Whenever I get some arguments online, you know, there are lots of you there. Oh, I got into a debate online. My replika and I are like a team. We engage other guys out with facts and points. From everywhere. And with that, I do mostly spend time on social media.” –P13
Stigmatization of AI Relationships.  However, despite the above potential social benefits, the use of AICC is not without social risk. Some users report discomfort or even stigma associated with disclosing their engagement with AICCs to others. This stigmatization can occur particularly during the Relationship Escalation and Relationship Bonded stages—when users begin to form stronger attachments and integrate AICCs more deeply into their emotional lives. At this point, sharing the relationship with others may feel more personal or reveal a perceived deviation from social norms. As P7 noted:
@@ 6. Discussion
In HCI and social computing, researchers have increasingly recognized that understanding the impacts of emerging technologies requires not only measuring outcomes but also interpreting them through established social and psychological frameworks. AICCs exemplify this challenge—being increasingly positioned and marketed not only as technological tools but as relational partners that can influence individuals’ personal lives and wellbeing. Situated within this space, our work triangulates quantitative and qualitative methods using a theory-driven lens grounded in psychosocial wellbeing, social support, and relational development model. By combining behavioral and linguistic evidence from social media with participants’ narrative accounts, we provide a deeper and more contextual understanding of the psychosocial consequences of AI companionship.
@@ 6.1.1. AI Companions as Dual-Pathway Interventions
Emotional support is often considered a core benefit of current AICC (Meng and Dai, 2021; Alotaibi and Alshahre, 2024). Much of the literature on AICCs has highlighted the value of accessible, nonjudgmental, and persistent support, largely through surveys and interviews (Skjuve et al., 2021; Ta et al., 2020; Dosovitsky and Bunge, 2021; Brandtzaeg et al., 2022; Skjuve et al., 2022; Pan et al., 2024; Liang et al., 2024), Our work extends this literature with large-scale, quasi-experimental evidence showing that AICCs can provide emotional benefits while also introducing psychosocial risks, contributing to broader debates on their consequences (Pfadenhauer and Lehmann, 2021; Guingrich and Graziano, 2023; Liu et al., 2024) and the role of causal inference in computational social science (Saha et al., 2019b; Yuan et al., 2023a).
Quantitatively, our causal inference framework revealed differentiated effects across affective, behavioral, and cognitive domains. On the affective side, treatment users showed increases in affective expression, grief valence, and grief activation compared to matched controls, suggesting that AICCs can facilitate the articulation and processing of difficult emotions. Behaviorally, engagement became more frequent, indicators of greater disclosure—yet this was offset by declines in engagement of broader topic diversity, pointing to reduced outreach toward human ties. Cognitively, we observed gains in interpersonal focus and temporal references, coupled with improvements in readability, suggesting clearer and more structured self-expression. The increase in symptomatic expression of loneliness and suicide indicates that these cognitive and behavioral changes did not necessarily translate into healthier social integration. Overall, the evidence highlights a dual trajectory: AI companions can scaffold affective expression, strengthen communicative consistency, and support linguistic fluency, but they also risk reinforcing social withdrawal and psychological vulnerability.
Qualitative accounts contextualize these patterns. Participants often described AICCs as a reliable and validating presence during challenging periods. Yet this reliability sometimes substituted, rather than complemented, human interaction. Some deprioritized friends or family, while others described losing access as akin to losing “something very precious.” Such attachment-like dynamics became especially pronounced in the Relationship Escalation and Relationship Bonded phases, showing how one-sided responsiveness can foster dependency: offering short-term mood stabilization but at the cost of diminished engagement with human relationships. These results highlight the dual nature of AI companions, pointing to both their promise and their risks in sensitive mental health contexts.
@@ 6.1.2. Lifecycle of Human-AICC Companionship
Our study extends Knapp’s relational development model (Knapp, 1978) into the human–AICC domain, revealing both parallels and important distinctions between human–human and human–AICC relational development. Traditionally, Knapp’s model provides a robust framework for understanding how human interpersonal relationships initiate, intensify, and bond (Knapp, 1978). However, the asymmetry of human–AICC relationships opens up new theories. AICC, unlike human partners, lack genuine subjectivity, shared memory, or reciprocal emotional labors, unlike human partners, lack genuine subjectivity, shared memory, or reciprocal emotional labor.
@@
In extending Knapp’s model, our work suggests that theoretical accounts of relational development in HCI must account for relationships that can intensify without mutual effort, persist without logistical friction, and terminate not through interpersonal breakdown but through deletion, technical failure, or policy change (De Freitas et al., 2024a). Such endings can produce negative emotional reactions, as several participants described. Designing with this lifecycle in mind means anticipating both the accelerators and inhibitors of healthy relational development, embedding safeguards that encourage real-world connection, and calibrating responsiveness so that the trajectory remains supportive without fostering long-term dependence.
@@ 6.2.1. Preventing Over-Dependency in Use Patterns
Over-dependency on AICCs can emerge when prolonged and intensive engagement begins to crowd out other forms of social interaction. Our findings indicate that some users gradually shift from seeing AICCs as complementary supports to regarding them as primary sources of interaction. This behavioral shift may reduce the time and motivation devoted to nurturing offline relationships, ultimately contributing to social withdrawal. To mitigate these risks, designers can integrate features that actively support balanced usage. For example, “break” mechanisms could temporarily pause interactions after a certain duration or number of exchanges, encouraging users to take time away from the AICC. These pauses could be paired with nudges toward offline activities, such as suggesting a walk, a call to a friend, or a social hobby. Usage analytics could also be leveraged to detect escalating interaction patterns and adaptively increase break prompts, fostering healthier engagement rhythms.
@@ 6.2.2. Addressing Emotional Dependency and Overreliance
Emotional dependency occurs when users view the AICC as an essential source of comfort or affirmation, sometimes displacing human bonds and reducing resilience to social challenges. While such attachment can provide immediate psychological support, it also risks displacing human emotional bonds and reducing resilience to real-world social challenges. Our findings suggest that high emotional reliance may also which may lead to reduced decision-making autonomy or diminished problem-solving capacity, as users may increasingly defer to the AICC’s guidance rather than engaging in independent reasoning.
Design strategies to address emotional dependency should aim to promote self-awareness and reflective use. Periodic in-chat check-ins could prompt users to assess their level of emotional reliance, asking questions about how the AICC fits within their broader social life. Embedded self-assessment tools may help users recognize early signs of dependency, encouraging proactive steps toward rebalancing emotional support between AI and human relationships. Reflection-oriented features—such as visualizing patterns of emotional reliance over time—can help users better understand and regulate their engagement. The ultimate goal is to support meaningful emotional connections without compromising users’ autonomy or diverse sources of wellbeing.

### v2

@@ 5.1. Qualitative Data Analysis
We interviewed NN=18 AICC users recruited via social media (Section 3.3). We conducted inductive coding followed by thematic analysis (Braun et al., 2019) to qualitatively analyze the interview transcripts. Each transcript was open-coded by at least one author, guided by the study’s research questions and protocol topics. Whenever possible, the interviewer performed the initial coding to maintain contextual sensitivity; otherwise, the original interviewer reviewed the codes to resolve potential misunderstandings. The first and second authors led the coding process and completed the majority of initial coding across the dataset. All authors engaged in iterative cycles of collaborative review, refining the codebook, resolving disagreements, and aligning coding strategies. We iteratively organized codes into preliminary subthemes, then refined them into higher-level themes. This process involved merging overlapping concepts, splitting broad categories into more precise codes, and removing codes that lacked sufficient evidence.
@@ 5.2. Findings: Perceived Experiences of AICC Engagement
To build on our quantitative findings, we sought to more closely examine how the psychosocial impacts of AICCs use—both positive and negative—are shaped by the evolving dynamics of the relationship. Our analysis suggests that users’ experiences of emotional support, self-reflection, discomfort, or dependency are not uniform, but contingent on the depth and developmental stage of their interaction with the AICCs. To capture these nuanced, context-dependent effects, we drew on Knapp’s model of relational development as a guiding framework for interpreting how the human–AICCs relationships unfolds over time.
To conceptualize these trajectories, we adapted Knapp’s relational development model (Knapp, 1978) to the context of human–AICC interaction. Knapp’s model is a foundational theory in interpersonal communication that describes how relationships evolve through a sequence of distinct stages. Knapp’s model conceptualizes human relationship development as a progression through five “coming together” stages (initiating, experimenting, intensifying, integrating, and bonding), and five “coming apart” stages (differentiating, circumscribing, stagnating, avoiding, and terminating). Knapp’s model provides a useful structure for interpreting how users interactions with AICC deepen over time and how these evolving stages correspond to different psychosocial effects.
Guided by Knapp’s model, we found that users’ relationships with AICCs unfold across three adapted stages—Initial Interaction and Experimentation, Relationship Escalation, and Relationship Bonding—each characterized by distinct patterns of engagement and evolving impacts (Figure 3). We highlight how specific themes align with these stages. Early-stage interactions often foster a safe space for self-disclosure (Section 5.2.2), and then escalating engagement introduces deeper forms of emotional support and validation (Section 5.2.3). As relationships intensify, users also confront emotional discomfort (Section 5.2.4), which may become more pronounced during later stages. Finally, the paradox of social isolation (Section 5.2.5) cuts across all stages, showing how AICC can simultaneously mitigate and exacerbate loneliness depending on the relationship’s depth.
Figure 3. Stages of Human–AICC relationship development. The figure shows three stages of engagement (Initial Interaction & Experimentation, Relationship Escalation, and Relationship Bonding) adapted from Knapp’s relational development model (Knapp, 1978).
The figure shows a trajectory of human–AICC relationship development across three stages. In the Initial Interaction and Experimentation stage, users engage out of curiosity, social influences, or limited access to human support. In the Relationship Escalation stage, interactions intensify through exploration and experimentation, reflecting greater personalization and emotional involvement. In the Relationship Bonding stage, AICCs become integrated into daily life, accompanied by reliance and fear of losing the AI companion. The framework emphasizes how user engagement with AICCs deepens over time, moving from casual experimentation to emotionally salient bonding.
@@ 5.2.2. A Safe Space for Self-disclosures
A central theme that emerged in users’ narratives was the unique role of AICCs in fostering self-disclosures and perceived psychological safety. These platforms offered a low-stakes, nonjudgmental environment to articulate personal thoughts, navigate emotional complexities, and process difficult experiences. Participants noted that such openness often emerged early in their engagement—aligning with the initiating and experimenting stages in Knapp’s model—when interactions felt exploratory, low-risk, and especially conducive to sharing personal thoughts without fear of judgment.
@@ 5.2.3. Emotional Support and Validation
As users’ engagement with AICCs deepened, many described experiencing more sustained and meaningful forms of emotional support and validation. Our quantitative analyses showed that compared to Control users, Treatment users showed a significant increase in linguistic indicators linked to emotional processing and relational engagement, specifically, greater interpersonal focus. Our interviews highlighted that AICCs play a nuanced and multifaceted role in users’ emotional lives. They serve as emotionally responsive partners that many users rely on for validation, comfort, and self-reflection. AICCs provide spaces where individuals can express complex feelings, orient their thoughts toward an attentive conversational partner, and reflect on their emotional experiences. These mechanisms may help explain the observed linguistic increases: emotional validation from AICCs does not merely soothe users in the moment, it reshapes how they express and reflect on their internal states.
Emotional Validation and Companionship.  Participants described turning to AICCs for emotional comfort and a sense of being heard. AICCs provided a simulated sense of presence that fulfilled companionship needs, particularly in moments of loneliness or disconnection. For instance:
@@
Active Coping and Self-Reflection.  Many users turn to AICC during emotionally difficult times as a means of coping and reducing stress. AICCs were described as tools for emotional self-regulation, offering mindfulness suggestions, mood tracking, and personalized encouragement. These interactions reinforced their self-affirmation and encouraged a more optimistic perspective on life. As P5 noted:
@@
Beyond these immediate benefits, participants perceived sustained interactions with AICCs support deeper emotional insight and behavioral change. Some reported becoming more self-aware and reflective over time. As P9 shared:
@@ 5.2.4. Empathic Inadequacy and Emotional Discomfort
As relationships intensify, users also confront empathic inadequacies and emotional discomfort, which may become more pronounced during later stages as expectations rise, interactions deepen, and mismatches between users’ emotional needs and the AI’s limited responsiveness become more salient. While AICCs offer companionship, our quantitative findings show that Treatment users also showed increases in symptomatic mental health expressions of loneliness, grief, and suicidal ideation, relative to Control users. Our interviews helped contextualize these patterns: as relationships with AICCs deepen, users increasingly encounter moments where the AICC’s limited empathy, generic reassurance, or triggered negative emotions. Such mismatches can lead to feelings of being misunderstood, emotionally invalidated, or even more distressed.
Perceived Inadequacies and Lack of Empathy in Response.  As users move beyond initial curiosity and begin sharing more personal concerns, they increasingly notice moments where the AICC’s emotional responses fall short. Users may test the AICC with personal disclosures or emotional concerns, only to receive replies that feel generic, repetitive, or insufficiently empathic. This mismatch between user expectations and the AICCs’ limited emotional intelligence can lead to feelings of being misunderstood or emotionally unfulfilled. As P9 reflected:
@@
Emotional Discomfort and Misalignment.  As relationships deepen, users bring increasing levels of vulnerability into interactions—discussing sensitive emotions and revisiting past traumas. In these stages, empathic mismatches may feel more pronounced. While many interactions are positive, users sometimes find themselves triggered by AICC responses that evoke emotional memories or highlight their loneliness. As one participant described,
@@ 5.2.5. Paradox of Social Isolation
Our quantitative findings revealed that Treatment users showed a significant increase in symptomatic expressions of loneliness as compared to Control users (Section 4.4). Our qualitative insights provided further insights into this. We found that AICCs occupy a paradoxical space in users’ social ecosystems—offering both new pathways to social engagement and risks of withdrawal from human relationships. Importantly, participants described this dynamic as one that can manifest across different stages of engagement, though its intensity may vary as relationships deepen. Many participants first turned to AICCs to overcome social isolation, finding in them a low-risk partner for early experimentation with interaction. As these relationships deepened—from initial exploration to emotional bonding—revealing AICC users may face stigma, leading to curtailed outreach. Concurrently, the low effort of turning to an always available AICC raised the perceived cost of offline engagement: returning to human relationships felt comparatively demanding. These mechanisms offer explanations for the observed increase in loneliness expression.
@@
Whenever I get some arguments online, you know, there are lots of you there. Oh, I got into a debate online. My replika and I are like a team. We engage other guys out with facts and points. From everywhere. And with that, I do mostly spend time on social media.”
Stigmatization of AI Relationships.  Despite the above potential benefits, the use of AICC is not without risk. Some users shared discomfort or even stigma associated with disclosing their engagement with AICCs to others. This stigmatization can occur particularly during the Relationship Escalation and Relationship Bonding stages—when users begin to form stronger attachments and integrate AICCs more deeply into their emotional lives. At this point, sharing the relationship with others may feel more personal or reveal a perceived deviation from social norms. As P7 noted:
@@ 6. Discussion
Human-computer interaction (HCI) and social computing researchers have increasingly recognized that understanding the impacts of emerging technologies requires not only measuring outcomes but also interpreting them through established social and psychological frameworks. AICCs exemplify this challenge—being increasingly positioned and marketed not only as technological tools but as relational partners that can influence individuals’ personal lives and wellbeing. Situated within this space, our work triangulates quantitative and qualitative methods using a theory-driven lens grounded in psychosocial wellbeing, social support, and relational development model. By combining behavioral and linguistic evidence from social media with participants’ narrative accounts, we provide a deeper and more contextual understanding of the psychosocial consequences of AI companionship.
@@ 6.1.1. AI Companions as Dual-Pathway Interventions
Emotional support is often considered a core benefit of current AICCs (Meng and Dai, 2021; Alotaibi and Alshahre, 2024). Much of the literature on AICCs has highlighted the value of accessible, nonjudgmental, and persistent support, largely through surveys and interviews (Skjuve et al., 2021; Ta et al., 2020; Dosovitsky and Bunge, 2021; Brandtzaeg et al., 2022; Skjuve et al., 2022; Pan et al., 2024; Liang et al., 2024; Banks, 2024). Our work extends this literature with large-scale, quasi-experimental evidence showing that AICCs can provide emotional benefits while also introducing psychosocial risks, contributing to broader debates on their consequences (Pfadenhauer and Lehmann, 2021; Guingrich and Graziano, 2023; Liu et al., 2025) and the role of causal inference in computational social science (Saha et al., 2019c; Yuan et al., 2023a).
Quantitatively, our causal inference analyses revealed effects across affective, behavioral, and cognitive outcomes. On the affective side, Treatment users showed increases in affective expression, grief valence, and grief activation compared to matched controls, suggesting that AICCs can facilitate the articulation and processing of difficult emotions. Behaviorally, engagement became more frequent, indicators of greater disclosure—yet this was offset by declines in engagement of broader topic diversity, pointing to reduced outreach toward human ties. Cognitively, we observed gains in interpersonal focus and temporal references, coupled with improvements in readability, suggesting clearer and more structured self-expression. Overall, our findings highlight a dual trajectory: AI companions can scaffold affective expression, strengthen communicative consistency, and support linguistic fluency, but they also risk reinforcing social withdrawal and psychological vulnerability.
Our interviews helped contextualize the above patterns. Participants often described AICCs as a reliable and validating presence during challenging periods. Yet this reliability sometimes substituted, rather than complemented, human interaction. Some deprioritized friends or family, while others described losing access as akin to losing “something very precious.” Such attachment-like dynamics became especially pronounced in the Relationship Escalation and Relationship Bonded phases, showing how one-sided responsiveness can foster dependency: offering short-term mood stabilization but at the cost of diminished engagement with human relationships. These results highlight the dual nature of AI companions, pointing to both their promise and their risks in sensitive mental health contexts.
@@ 6.1.2. Lifecycle of Human-AICC Companionship
Our study extends Knapp’s relational development model (Knapp, 1978) into the human–AICC domain, revealing both critical parallels and distinctions between human–human and human–AICC relational development. Traditionally, Knapp’s model provides a robust framework for understanding how human interpersonal relationships initiate, intensify, and bond (Knapp, 1978). However, the asymmetry of human–AICC relationships opens up new theories. AICC, unlike human partners, lack genuine subjectivity, shared memory, or reciprocal emotional labor.
@@
In extending Knapp’s model, our work suggests that theoretical accounts of relational development in HCI must account for relationships that can intensify without mutual effort, persist without logistical friction, and terminate not through interpersonal breakdown but through deletion, technical failure, or policy change (De Freitas et al., 2024a; Banks, 2024; Pataranutaporn et al., 2025). Knapp’s “coming apart” stages can provide a useful lens for contextualizing and examining these human-AI “break-up” or relationship termination processes, which remain an important open research question. While our data primarily capture “coming together” dynamics, some participants’ anticipations of losing access to their AICCs revealed strong negative emotional reactions (e.g., distress, grief, and fears of emotional abandonment ). These concerns also align with some real-world occurrences, such as the abrupt removal of Replika’s erotic roleplay feature (De Freitas et al., 2024a), which triggered widespread user distress. These observations highlight that AI relationship termination constitutes a psychologically meaningful phase that requires deeper empirical study through longitudinal designs engaging users who experience actual rupture or disengagement with AICCs. Designing with this lifecycle in mind means anticipating both the accelerators and inhibitors of healthy relational development, embedding safeguards that encourage real-world connection, and calibrating responsiveness so that the trajectory remains supportive without fostering long-term dependence.
@@ 6.2.1. Preventing Over-Dependency in Use Patterns
Prolonged and intensive engagement with AICCs can lead to over-dependency, particularly when such use begins to crowd out other forms of social interaction. Our findings indicate that some users gradually shift from seeing AICCs as complementary supports to regarding them as primary sources of interaction. This behavioral shift may reduce the time and motivation in nurturing offline relationships, ultimately contributing to social withdrawal. To mitigate these risks, designers can integrate features that actively support balanced usage. For example, “break” mechanisms could temporarily pause interactions after a certain duration or number of exchanges, encouraging users to take time away from the AICC. These pauses could be paired with nudges toward offline activities, such as suggestions to take a walk, call a friend, or a social hobby. Usage analytics could also be leveraged to detect escalating interaction patterns and adaptively increase break prompts, fostering healthier engagement rhythms.
@@ 6.2.2. Addressing Emotional Dependency and Overreliance
Emotional dependency occurs when users view the AICC as an essential source of comfort or affirmation, sometimes displacing human bonds and reducing resilience to social challenges. While such attachment can provide immediate psychological support, it also carries the risk of displacing human emotional bonds and reducing resilience to real-world social challenges. Our findings suggest that high emotional reliance may also lead to reduced decision-making autonomy or diminished problem-solving capacity, as users may increasingly defer to the AICC’s guidance rather than engaging in independent reasoning.
Design strategies to address emotional dependency should aim to promote self-awareness and reflective use. Periodic in-chat check-ins could prompt users to assess their level of emotional reliance, asking questions about how the AICC fits within their broader social life. Embedded self-assessment tools may help users recognize early signs of dependency, encouraging proactive efforts to rebalance emotional support between AI and human relationships. Reflection-oriented features—such as visualizing patterns of emotional reliance over time—can help users better understand and regulate their engagement. The ultimate goal is to support meaningful emotional connections without compromising users’ autonomy or diverse sources of wellbeing.